The soln of the dfferential eqn   2x dy/dx - y  = 3 represents


(a) straight line      (b) circle    (c) parabola       (d) ellipse


 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit

2x dy/dx - y  = 3

2xdy/dx =y+3

2 dy/(y+3) =dx /x

2ln(y+3) =ln x +lnc

(y+3)2 =cx

this is equation of parabola


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin